# A Coalition Game for On-demand Multi-modal 3D Automated Delivery System

### v1

@@ 3 Problem Description
3.1 Setup and Stylized Case Study
In this study, we adopt the transportation network of Mississauga, Canada (Figure 1(a)) to simulate the pickup and delivery problem with time windows as a food delivery with multiple depots and agents. We obtain data from the OpenStreetMap city map using OSMnx (Boeing, 2017) as the region of interest is shown in Figure 1(a), which includes a restricted area due to the presence of the airport near the case study. This network represents the ground on which ADRs and UAVs operate. The street’s intersections denote the nodes connected by corresponding edges. Each edge poses certain transportation features, such as constant path length and real-time traffic speed, resulting in the dynamic travel time associated with each link obtained from the OSMnx Python package. Restaurants as destinations where vehicles would be assigned to pick up food are collected utilizing the web-based data mining tool for Open Street Map, overpass turbine, and the footprint and height of the buildings in the area (Figure 1(b)). In addition, the demand used in this study is time-dependent based on the time of order during the evening peak and the size of the network after (Samouh et al., 2020).
(a) The network zone boundary.
(b) Mississauga city transportation network.
Figure 1: Case study network.
3.2 Cooperative Multi-Modal On-Demand Delivery Problem
Before delving into how the system is optimized to operate sustainably in the urban environment, we must first define how vehicles operate. The ADR traverse along any link except those leading to highways, like primary and secondary roads. This stems from the municipal and safety regulations and operational limitations due to battery constraints. Therefore, the only region allocated to the ADRs is designated residential areas where the ADR can move along sidewalks. The residential edges are in blue in Figure 2, which is part of the case study area better to illustrate the network. On the other hand, the notion of the road network is adopted for aerial delivery. Primary and secondary roads are utilized as the main delivery routes for UAVs, which extend at multiple levels in the airspace with a limit of operation of 400 feet according to the (Osler, 2021). This has the benefit of moving in the structured paths above the roads to bypass the high rises and city infrastructure, as can be visible in Figure 3, which shows a green route for direct delivery or a red path for the next shortest route delivery due to obstruction. The UAV will freely move horizontally and vertically along the tubes and traverse straight in the absence of obstacles. In this study, optimized performance is considered a move in space during any delivery and take-off and landing occur from a constant level. As a result, the aerial network replicates the double-lane ground roads with the capability of altitude transition at primary intersection nodes. If municipal regulations permit, UAVs can travel through straight lines when the airspace is clear. Otherwise, they move through the edges of the elevated road network as depicted in Figure 3. Similar to ADRs, UAVs pose some limitations in certain areas. For example, the red arced region in Figure 3 shows the restricted circular area near Pearson airport, requiring at least a $5.6\;km$ distance.
Figure 2: UAV Network Routing
Figure 3: Aerial network for UAVs including restricted zone marked as red arced region
4 Methodology
This section describes how a capacitated pick-up and delivery problem with time windows (CPDPTW) is defined, followed by our solution approach for cost optimization and cooperative strategy. We first present the notation and problem formulation using mathematical expressions as well as a learning-based approach. Next, a vehicle routing game is defined to obtain multi-modal coalition operational costs. The validity of the core properties of coalition theory is confirmed in or obtains an optimal cooperation strategy.
@@ 4 Methodology
Initially, the delivery problem network comprises two unique graphs for a fleet of UAVs and ADRs based on the road network of an urban area. The nodes are represented by a directed graph for UAV network $G^{d}=(P\cup D,E^{d})$, where $P=\left\{x_{1},\ldots,x_{N}\right\}$ denotes the set of $N$ pickup nodes, and $D=\left\{x_{N+1},\ldots,x_{2N}\right\}$ as corresponding delivery nodes. In addition, $E^{d}=$$\left\{(i,j)\mid x_{i},x_{j}\in X\right\}$ denotes the set of edges connecting the locations. The same representation is applied for the ADR graph network $G^{r}=(P\cup D,E^{r})$ with the same node but different edges. The coordinate of the $i$th location, order demand and time window is denoted by $\mathbf{u}_{i}$, $q_{i}$, and $\left[e_{i},l_{i}\right]$ respectively with $q_{i}>0$ and $q_{i+N}=-q_{i}$, for corresponding delivery node. Each vehicle must serve the pickup and delivery of requests together, accounting for precedence constraints within the time window of each point; otherwise, they get delayed, and the penalty is considered. There are $N^{d}$ and $N^{r}$ UAVs and ADRs correspondingly. The $k$th vehicle, $k\in\left\{1,\ldots,N^{d,r}\right\}$, has a capacity $Q_{k}$ and a battery size $B_{k}$. If the $k$th vehicle is used, it departs from a depot with a sequence of locations and returns to a depot after its final delivery or when it needs recharging. In other words, a delivery scenario has $n$ customer requests, and each is constituted into a pickup task $i\in P$ and a delivery task $n+N\in D$. A vehicle with a fully charged battery is sent to accomplish all the assigned tasks; the vehicle can recharged at any depot $c\in C$.
@@
Furthermore, the travel time between any two nodes is the time on the road network based on the impedance weight computed by Dijkstra’s algorithm. Note that the vehicle’s graph is built upon the customer nodes (pickup and delivery), and every other node is considered an intermediary point that connects these nodes, at which vehicles would not stop. The UAV and ADR networks vary by wind and congestion, respectively, meaning their impedance updates at different times of the day. We adopted the energy consumption model for UAVs in the effect of the wind as proposed in (Liu et al., 2023). Solving the vehicle delivery problem aims to find a route that minimizes the delivery time, including the delay respected to time windows (Santiyuda et al., 2024), while being subjected to constraints regarding the delivery mission and vehicle properties. Further assumptions used in this study are listed as follows:
@@
    Both modes can be assigned for more than one order.
@@
    Each mode has a certain battery consumption and capacity as well as the battery lower bound threshold to return to the nearest depot for recharging.
@@
The mathematical notations and formulation of the Electric Pickup and Delivery Problem with Time Windows (E-PDPTW) are proposed as in A and Table 5 depicts the variables and parameters definition in this notation. In what follows, the problem model and dynamic are defined by the Markov decision process.
@@ 4.1 Markov Decision Process
A generic MDP consists of four components: state space $S$, action space $A$, reward function $r(s,a)$, and state transition probability $p\left(s^{\prime}\mid s,a\right)$. Specifically, for the PDPTW described earlier in this paper, these components are defined as follows:
@@
    States: Composed of the graph vertex state and the vehicle state, denoted as $s_{t}={\left\{x_{t},v_{t}\right\}}$ at step t, where $x_{t}=(u,q_{t},e_{t},l_{t})$. The kth vehicle state $v^{k}_{t}$ is composed of its load $u_{t}$, battery level $e_{t}$, and traveled time $\tau^{t}$, expressed as $v_{t}=\left[\tau^{k}_{t},u^{k}_{t},e^{k}_{t}\right]$.
@@
    Actions: $a_{t}$ determines the vehicle’s node selection at step t. The sequence of actions generated from the initial to the final step should be combinations of nodes starting and ending with the depots.
@@
    Reward: PDPTW aims to minimize fleet delay and travel time. Therefore, the reward function is given in Equation 1.
    $\;R=\alpha_{1}\sum_{k\in N^{d}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\alpha_{2}\sum_{k\in N^{r}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\sum_{i\in P\cup D}\alpha_{3}\left|T_{ik}-e_{i}\right|+\sum_{i\in P\cup D}max\left\{T_{ik}-l_{i},0\right\}$ (1)
    Another term, Equation 2, is added to the reward when the agent’s energy falls below the battery threshold as a penalty penalizes the agent whose battery has run out. Ensuring efficient behaviour and battery usage encourages the agents to avoid poor routing decisions, leading to increased operational costs and low-battery situations.
@@
    Where $\lambda$ is a positive coefficient, and $R^{t}_{k}$ is the reward of the agent $k$ at step $t$.
@@
    Transition: The system state will be updated from $S_{t}$ to $S_{t+1}$ based on the currently executed action $a_{t}$. The dynamic features of the problem, such as vehicle load, battery level, and travelled time, are being changed through consecutive nodes based on the vehicle’s features (Equations 3, 4, 5, and 6). First, the system time is updated based on these equations.
    $\tau^{t+1}=\left\{\begin{array}[]{l}\max\left(\tau^{t},l_{i}\right)+s+t_{ijk},\text{ if }i\in P\cup D\\ \tau^{t}+\left(B^{k}-e^{k}_{t}\right)/\eta_{k}+t_{ijk},\text{ if }i\in C\end{array}\right.$ (3)
    where $\eta_{k}$ is the charging rate to charge the battery from the given level for any vehicle $k$, $s$ is a constant representing the service time at each customer vertex. Next, the battery level of the vehicle is updated:
@@
    Where $e_{ijk}$ is the energy consumption of the vehicle $k$ travelling from vertex $i$ to vertex $j$. The power consumption model for UAV and ADR batteries is discussed in B.
    Finally, the vehicles load $u^{t}$, and the remaining demand, $d_{i}^{t}$, at each vertex are updated as follows.
@@
    Policy: The stochastic policy $p_{\theta}$ automatically selects a node at each time step under the precedence constraint. This process is repeated iteratively until all pickup-delivery services are completed. The final outcome engendered by performing the policy is a permutation of all nodes, which prescribes the order of each node for the vehicle to visit, i.e., $\pi=\left\{\pi_{0},\pi_{1},\ldots,\pi_{T}\right\}$. Based on the chain rule, the probability of an output solution is factorized as Equation 7.
    $P(\pi\mid X)=\prod_{t=0}^{T-1}p_{\theta}\left(\pi_{t}\mid X,\pi_{1:t-1}\right)$ (7)
    where $X$ is the input of a problem instance. The decision-making about the node selection will be performed based on the learned $p_{\theta}$.
Afterwards, the detailed architecture of the transformer and elements of the encoder and decoder are elaborated.
4.2 Reinforcement Learning Approach
This study integrates a Graph Attention Network (GAT) and Transformer architecture to address these challenges using an end-to-end learning approach. Deep reinforcement learning (DRL), augmented with a dynamic transformer model and a novel customized graph attention network, solves PDPTW and minimizes the delivery time delay. The proposed approach coordinates a multi-modal fleet with a network on the ground and in the air, using a dual encoder architecture to capture the embedding of both modes, along with heterogeneous attention to account for precedence constraints (Zhang et al., 2022) and, more importantly, a customer’s spatial and temporal correlation to enrich node and graph context embedding for dynamic routing. A dynamic encoding mechanism is introduced to update edges in addition to node embedding, serving as a dynamic encoder that reflects the online routing problem and addresses on-demand delivery. The flowchart of this methodology is shown in the Figure 4.
Figure 4: The process of routing problem solution

### v2

@@ 3 Problem Description
The research problem is to minimize the vehicles’ travel time and customer waiting time for on-demand last-mile delivery in an urban environment using two distinct sets of autonomous vehicles. To be more specific, a time-sensitive scenario, such as food delivery in peak hour demand at a city network level, is considered to be solved for the given multi-modal fleet. In what follows, the preliminaries and mathematical formulation are provided, followed by a discussion on a complete operational network in the case study for UAV and ADR.
@@
Initially, the delivery problem network comprises two unique graphs for a fleet of UAVs and ADRs based on the road network of an urban area. The nodes are represented by a directed graph for UAV network $G^{d}=(D^{\prime}\cup P\cup D,E^{d})$, where $P=\left\{x_{1},\ldots,x_{n}\right\}$ denotes the set of $n$ pickup nodes, $D=\left\{x_{n+1},\ldots,x_{2n}\right\}$ as corresponding delivery nodes, and $D^{\prime}=\left\{x_{0},\ldots,x_{|D^{\prime}|}\right\}$ are depots set. $x_{i}$ denotes the location of node $i$, and $|D^{\prime}|$ shows the number of depots. In addition, $E^{d}=$$\left\{e^{d}_{ij}\mid(i,j)\in N\right\}$ denotes the set of edges connecting the locations and $N=\left\{P\cup D\cup D^{\prime}\right\}$ represent all the nodes in the graph. The same representation is applied for the ADR graph network $G^{r}=(D^{\prime}\cup P\cup D,E^{r})$ with the same nodes but different edges $E^{r}=$$\left\{e^{r}_{ij}\mid(i,j)\in N\right\}$. The $i$th order demand and time window is denoted by $q_{i}$ and $\left[e_{i},l_{i}\right]$ ($e_{i}$ for pickup and $l_{i}$ for delivery node), respectively with $q_{i}>0$ and $q_{i+N}=-q_{i}$, for corresponding delivery node. Each vehicle must serve the pickup and delivery of requests together, accounting for precedence constraints within the time window of each point; otherwise, they get delayed, and the penalty is considered. There are $N^{d}$ and $N^{r}$ UAVs and ADRs, respectively. The $k$th vehicle, $k\in\left\{1,\ldots,K\right\}$, where $K$ is total number of vehicles, has a capacity $Q_{k}$ and a battery size $B_{k}$. If the $k$th vehicle is used, it departs from a depot with a sequence of locations and returns to a depot after its final delivery or when it needs recharging. In other words, a delivery scenario has $n$ customer requests, and each is constituted into a pickup task $i\in P$ and a delivery task $i+n\in D$. A vehicle with a fully charged battery is sent to accomplish all the assigned tasks; the vehicle can be recharged at any depot.
@@
Furthermore, the initial graph is sampled from a larger graph that is generated randomly with a number of nodes five times greater than the size of pickup and delivery nodes, and edges are also generated randomly from half of the maximum possible number of edges to full size. By doing so, after sampling the delivery network graph, the nodes’ locations remain the same, although the edges are not necessarily a direct link. In this regard, the travel time between any two nodes is the time on the road network based on the impedance weight computed by Dijkstra’s algorithm. Note that the vehicle’s graph is built upon the customer nodes (pickup and delivery), and every other node is considered an intermediary point that connects these nodes, at which vehicles would not stop. The UAV network varies by wind, meaning its impedance updates at different weather conditions. We adopted the energy consumption model for UAVs in the effect of the wind as proposed in (Liu et al., 2023). Solving the vehicle delivery problem aims to find a route that minimizes the delivery time, including the delay respected to time windows (Santiyuda et al., 2024), while being subjected to constraints regarding the delivery mission and vehicle properties. Further assumptions used in this study are listed as follows:
@@
    Both modes can serve more than one request simultaneously.
@@
    Each mode has a certain battery consumption and capacity, as well as the battery lower bound threshold to return to the nearest depot for recharging.
@@ 3.1 Mathematical Formulation
The mathematical notations and formulation of the Electric Pickup and Delivery Problem with Time Windows (CE-PDPTW) are proposed as follows, and Table 2 depicts the variables and parameters definition in this notation. The objectives and constraints in the mathematical formulation are listed as follows. The objective function is shown in Equation 5.
$\text{min}\;\alpha_{1}\sum_{k\in N^{d}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\alpha_{2}\sum_{k\in N^{r}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\sum_{i\in P\cup D}\alpha_{3}\left|T_{ik}-e_{i}\right|+\alpha_{4}\sum_{i\in P\cup D}max\left\{T_{ik}-l_{i},0\right\}$ (5)
Where $X_{ijk}$ is the binary variable equal to 1 if the vehicle $k$ travels from node $i$ to $j$. $t_{ijk}$ is time travelled by vehicle $k$ between node $i$ and $j$, $T_{ik}$ is the arrival time of the vehicle $k$ at node $i$, $\alpha_{1}$ and $\alpha_{2}$ are positive monetary conversion factors pf utilizing UAVs and ADRs respectively per hour. Also, the $\alpha_{3}$ and $\alpha_{4}$ are the monetary coefficients of waiting time penalty for customers that the delivery company has to compensate in the case of delayed delivery, in case of lateness at the pickup and delivery node, respectively. Additionally, the equations that constrain this objective function are discussed below.
$\sum_{j=0}^{P\cup D}X_{ijk}=1,k\in K\;\forall i\in D^{\prime}$ (6)
$\sum_{i=0}^{P\cup D}X_{ijk}=1,\forall k\in K\;\forall j\in D^{\prime}$ (7)
$\sum_{k=1}^{N^{r}\cup N^{d}}\sum_{i=0}^{N}X_{ijk}=1,\forall i\in\{P\cup D\}$ (8)
$\sum_{i=0}^{P\cup D}X_{ijk}=\sum_{i=0}^{P\cup D}X_{i\left(j+n\right)k},\forall j\in P,\;\forall k\in K$ (9)
$T_{ik}\geq e_{i},\forall k\in K,\forall i\in\{P\cup D\}$ (10)
$T_{ik}\leq T_{k,i+n},\forall i\in\left\{P\right\},\forall k\in K$ (11)
$T_{ki}+t_{ijk}+T_{r}^{k}R_{ik}\leq T_{kj}+(1-M)X_{ijk},\forall k\in K,\forall(i,j)\in\{P\cup D\}$ (12)
$0\leq u^{k}_{i}\leq Q_{k},\forall k\in K,\forall i\in\{P\cup D\}$ (13)
$u^{k}_{i}=u^{k}_{j}+q_{j}+(1-M)X_{ijk},\forall k\in K,\forall(i,j)\in\{P\cup D\}$ (14)
$e^{k}_{i}-e_{ijk}+B^{k}\left(1-X_{ijk}\right)\geq e_{j},\forall k\in K,\forall(i,j)\in\{P\cup D\}$ (15)
$e_{min}\leq e^{k}_{i}\leq B^{k},k\in K,\forall i\in\{P\cup D\}$ (16)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e_{ik}\geq e_{min}-M(1-R_{ik}),\;\forall i\in D^{\prime},\forall k\in K}$ (17)
$e^{k}_{i}=B^{k},\forall k\in K,i\in D^{\prime}$ (18)
$X_{ijk}\in\{0,1\},\forall(i,j)\in\{P\cup D\},\forall k\in K$ (19)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R_{ik}=0,\;\forall i\notin D^{\prime},\forall k\in K}$ (20)
The constraints in Equations 6 and 7 ensure the vehicles depart from and return to any depot, whether for recharging or ending the tour. The Equations 8 and 9 ensure the requests are only served once and by the same vehicle. The constraints in Equations 10, 11, 12 ensure that vehicles cannot start service at a location earlier than its early time window and cannot be later than the service at its corresponding delivery location, and $M$ is a large positive number. All vehicles must service the pickup location earlier than the delivery node. Besides, if the vehicle goes to the depot for recharging, it should wait at the depot to be recharged by $T_{r}^{k}$, which is the recharging time for vehicle $k$. The capacity constraint of the vehicles and the load balance update are given in Equations 13 and 14, with $u^{k}_{i}$ denoting the $k$th vehicle’s load after the service at the $i$th location is done. Equations 15 and 16 denote battery level update and upper and lower battery capacity constraints, respectively, where $e_{ijk}$ is the energy consumption of the vehicle $k$ travelling from node $i$ to node $j$. Equations 17 and 18 hold for recharging criteria and battery state after visiting depots, with $e^{k}_{i}$ denoting the $k$th vehicle’s battery level after going to the $i$th location. A binary variable of $R_{ik}$ is defined to track if vehicle $k$ recharges at node $i$. Besides, $e_{min}$ is the minimum battery threshold below which the vehicle battery level cannot be. Lastly, Equation 19 notes that the binary variable, and Equation 20 notes the recharging variable to be zero when not visiting a depot node.
Table 2: Parameters and Definitions
Parameter Definition
$D^{\prime}$ Depot instances set
$P$ Set of pickup nodes
$D$ Set of Delivery nodes
$N$ Set of all nodes
$E^{d}$ Set of UAV’s edges network
$E^{r}$ Set of ADR’s edges network
$[e_{i},l_{i}]$ Early and late time window
$x_{i}$ location of node $i$
$q_{i}$ Demand of customer $i$
$N^{d}$ Number of UAVs
$N^{r}$ Number of ADRs
$K$ Number of vehicles
$Q_{k}$ Capacity of vehicle $k$
$B_{k}$ Battery level of vehicle $k$
$n$ Number of customers
$T_{ik}$ Arrival time of the vehicle $k$ at node $i$
$T^{k}_{r}$ Recharging time for vehicle $k$
$t_{ijk}$ Time traveled by vehicle $k$ between node $i$ and $j$
$e_{ijk}$ Energy consumption of the vehicle $k$ travelling from node $i$ to node $j$.
$e_{min}$ The minimum battery capacity of vehicle $j$
$e^{k}_{i}$ battery level of vehicle $k$ at node $i$
$u^{k}_{i}$ Load of vehicle $k$ at node $i$
$X_{ijk}$ Binary decision variable for whether the vehicle $k$ travels from node $i$ to node $j$
$R_{ik}$ Binary variable, 1 if vehicle $k$ recharges at node $i$ (only allowed at depots)
3.2 Setup and Stylized Case Study
In this study, we adopt the transportation network of Mississauga, Canada (Figure 3(a)) to collect the essential information for data processing, such as nodes and edges of both ADR and UAV networks. We obtain data from the OpenStreetMap city map using OSMnx (Boeing, 2017) as the region of interest is shown in Figure 3(a), where the two-dimensional projection of the UAV network is shown. Furthermore, a more extensive network of the case study is depicted in Figure 3 as the ADR operation network bounded to the residential area and traversing along any link except those leading to highways, like primary and secondary roads. This stems from the municipal and safety regulations and operational limitations due to battery constraints. Therefore, the only region allocated to the ADRs is designated residential areas where the ADR can move along sidewalks. The residential edges are in yellow in Figure 4, which is part of the case study area, to better illustrate the network. Similar to ADRs, UAVs pose some limitations in certain areas. For example, the red arced region in Figure 1 shows the restricted circular area near Pearson airport, requiring at least a $5.6\;km$ distance. As a result, they interactively operate to complement their delivery tasks by joint environment coverage.
(a) The network zone boundary.
(b) Mississauga city transportation network.
Figure 3: Case study network.
On the other hand, the notion of the road network is adopted for aerial delivery. Primary and secondary roads are utilized as the main delivery routes for UAVs, which extend at multiple levels in the airspace with a limit of operation of 400 feet according to the (Osler, 2021). This has the benefit of bypassing the high rises and city infrastructure, as can be visible in Figure 3(a), which shows a blue route for direct delivery or a red path for the next shortest route delivery due to obstruction. The UAV will freely move horizontally and vertically along the tubes and traverse straight in the absence of obstacles. If municipal regulations permit, UAVs can travel through straight lines when the airspace is clear. Otherwise, they move through the edges of the elevated road network as depicted in Figure 1.
Figure 4: UAV Network Routing. This figure shows a top view of part of Mississauga with the pattern of the urban infrastructures labelled in yellow. The height heatmap is placed on the top right corner, which shows buildings’ height from ground level to the highest building in the network. In case of delivery within a cluttered space, the shortest path instead of the direct path is used.
Moreover, the street’s intersections, restaurants, and parking facilities denote the delivery, pickup, and depot nodes, respectively, connected by corresponding edges, where each node is obtained from the OSMnx Python package. The graph data, as well as the footprint and height of the buildings, are collected utilizing the web-based data mining tool for Open Street Map, overpass turbine. In addition, the demand order arrival time used in this study is based on the Poission distribution during the evening peak.
4 Methodology
This section describes our solution approach for cost optimization and the cooperative strategy of the CE-CPDPTW problem. We first present the reinforcement learning Markov decision process, followed by the encoder and decoder design and the multi-agent policy gradient training algorithm for sequential routing learning. Lastly, the coalition game application to the CE-CPDPTW is defined through a pipeline to check how agents can work more efficiently by forming a coalition by harnessing the core properties and the proposed cost allocation mechanism. In what follows, the problem model and dynamics are defined by the Markov decision process.
@@ 4.1 Markov Decision Process
A generic MDP consists of four components: state space $\mathcal{S}$, action space $\mathcal{A}$, reward function $r(s,a)$, and state transition probability $p\left(s^{\prime}\mid s,a\right)$. Each agent takes action in the environment and receives a signal reward and the probability of going to the next state, given its current action and state. Specifically, for the CE-CPDPTW environment described earlier in this paper, these components are defined as follows:
@@
    States: Composed of the graph node state and the vehicle state. The node state is $s_{t}=(x_{t},q_{t},e_{t},l_{t})$, consisting of location, demand, and time window of the node at time $t$. The kth vehicle state $v^{k}_{t}$ is composed of its load $u^{t}_{t}$, battery level $e^{k}_{t}$, and traveled time $\tau^{t}$, expressed as $v_{t}=\left[\tau^{k}_{t},u^{k}_{t},e^{k}_{t}\right]$ at a time $t$.
@@
    Actions: $a_{t}$ determines the vehicle’s node selection at step t. The sequence of actions generated from the initial to the final step should be a combination of nodes starting and ending with the depots.
@@
    Reward: This research aims to minimize fleet delay and travel time. Therefore, the reward function is given in Equation 21.
    $\;R=\alpha_{1}\sum_{k\in N^{d}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\alpha_{2}\sum_{k\in N^{r}}\sum_{(i,j)\in N}t_{ijk}X_{ijk}+\sum_{i\in P\cup D}\alpha_{3}\left|T_{ik}-e_{i}\right|+\alpha_{4}\sum_{i\in P\cup D}max\left\{T_{ik}-l_{i},0\right\}+r^{t}_{k}$ (21)
    Another term, Equation 22, is added to the reward when the agent’s energy falls below the battery threshold, penalizing the agent whose battery has run out. Ensuring efficient behaviour and battery usage encourages the agents to avoid poor routing decisions, leading to increased operational costs and low-battery situations.
@@
    Where $\lambda$ is a positive coefficient, and $R^{t}_{k}$ is the travel cost of agent $k$ at step $t$. Therefore, in case the battery exceeds the lower bound, $e_{min}$, a penalty proportional to the travel cost of visiting the unfinished customer is incurred.
@@
    Transition: The system state will be updated from $s_{t}$ to $s_{t+1}$ based on the currently executed action $a_{t}$. The dynamic features of the problem, such as vehicle load, battery level, and travelled time, are being changed through consecutive nodes based on the vehicle’s features (Equations 23, 24, 25, and 26). First, the system time is updated based on these equations.
    $\tau^{t+1}=\left\{\begin{array}[]{l}\max\left(\tau^{t},l_{i}\right)+t_{ijk},\text{ if }i\in P\cup D\\ \tau^{t}+\left(B^{k}-e^{k}_{t}\right)/\eta_{k}+t_{ijk},\text{ if }i\in C\end{array}\right.$ (23)
    where $\eta_{k}$ is the charging rate to charge the battery from the given level for any vehicle $k$. Next, the battery level of the vehicle is updated:
@@
    The initial battery size is given and updated based on the energy consumed in travelling from node $i$ to $j$, $e_{ijk}$. The power consumption model for UAV and ADR batteries is discussed in A. Their consumption model takes velocity and payload demand as input and computes the energy used in the battery.
    Finally, the vehicles load $u^{t}$, and the remaining demand, $d_{i}^{t}$, at each node are updated as follows.
@@
    Policy: The stochastic policy $p_{\theta}$ automatically selects a node at each time step. This process is repeated iteratively until all pickup-delivery services are completed with respect to the problem constraints. The final outcome engendered by performing the policy is a permutation of all nodes, which prescribes the order of each node for the vehicle to visit, i.e., $\pi=\left\{\pi_{0},\pi_{1},\ldots,\pi_{T}\right\}$. Based on the chain rule, the probability of an output solution is factorized as Equation 27.
    $P(\pi\mid s)=\prod_{t=0}^{T-1}p_{\theta}\left(\pi_{t}\mid s,\pi_{1:t-1}\right)$ (27)
    where $s$ is the state input to the problem instance. The decision-making about the node selection will be performed based on the learned $p_{\theta}$.
This policy is trained by a policy gradient algorithm, which is a multi-agent system learning the sequential task of the delivery problem by graph attention and a transformer model.
4.2 Reinforcement Learning Model
The overall framework of the RL model is to initially prepare the initial graph based on the problem states and each UAV and ADR network (aerial and terrestrial, as shown in Figure 2), as mentioned in the previous section. Subsequently, a MARL centralized controller is designed, comprising an encoder to extract the initial problem feature, and a decoder to learn the selection of nodes by contextual vehicle and routing information of the environment. Finally, an actor-critic training updates the policy parameters, leading to an optimal tour. Figure 5 demonstrates the problem methodology flowchart
Figure 5: Overview of the methodology
In the graph generation phase, node information is the same for both modes, like terrestrial for ADRs and aerial for UAVs; however, the edge features are different in terms of weight of connectivity, meaning that a node’s connections are the same, but with different weights. This Aerial network can be affected by urban building density. The more obstacles scattered along the customers’ location, the less direct a path can be built (refer to Figure 4; thereby, closer to the terrestrial network. These separate sets of edges, together with node features, are embedded by the dual encoder, which is discussed in the next section.